### v1

# Using Large Language Models to Tackle Fundamental Challenges in Graph Learning: A Comprehensive Survey

Abstract
Graphs are a widely used paradigm for representing non-Euclidean data, with applications ranging from social network analysis to biomolecular prediction. Conventional graph learning approaches typically rely on fixed structural assumptions or fully observed data, limiting their effectiveness in more complex, noisy, or evolving settings. Consequently, real-world graph data often violates the assumptions of traditional graph learning methods, in particular, it leads to four fundamental challenges: (1) Incompleteness, real-world graphs have missing nodes, edges, or attributes; (2) Imbalance, the distribution of the labels of nodes or edges and their structures for real-world graphs are highly skewed; (3) Cross-domain Heterogeneity, graphs from different domains exhibit incompatible feature spaces or structural patterns; and (4) Dynamic Instability, graphs evolve over time in unpredictable ways. Recent advances in Large Language Models (LLMs) offer the potential to tackle these challenges by leveraging rich semantic reasoning and external knowledge. This survey provides a comprehensive review of how LLMs can be integrated with graph learning to address the aforementioned challenges. For each challenge, we review both traditional solutions and modern LLM-driven approaches, highlighting how LLMs contribute unique advantages. Finally, we discuss open research questions and promising future directions in this emerging interdisciplinary field. To support further exploration, we have curated a repository of recent advances on graph learning challenges: https://github.com/limengran98/Awesome-Literature-Graph-Learning-Challenges.
keywords:
Graph Learning, Large Language Models, Graph Incompleteness, Data Imbalance, Cross-domain Graph Heterogeneity, Dynamic Graph Instability
[…]
[1]organization=Guangdong Key Laboratory of Intelligent Transportation System, School of Intelligent Systems Engineering, Shenzhen Campus of Sun Yat-sen University, city=Shenzhen, postcode=518107, state=Guangdong, country=China
[…]
[2]organization=University of Amsterdam, city=Amsterdam, country=The Netherlands
[…]
[3]organization=Aarhus University, city=Aarhus, country=Denmark
[…]
[4]organization=Beijing Institute of Artificial Intelligence, Beijing University of Technology, city=Beijing, postcode=100124, country=China
[…]
[5]organization=School of Information and Engineering, Chang’an University, city=Xi’an, postcode=710064, state=Shaanxi, country=China
[…]
[6]organization=Key Laboratory of Road and Traffic Engineering of the Ministry of Education, Tongji University, city=Shanghai, postcode=201804, country=China
[…]
[7]organization=Key Laboratory of Intelligent Transportation Technology and System, School of Transportation Science and Engineering, Beihang University, city=Beijing, postcode=100191, country=China
[…]
[8]organization=Jiangsu Key Laboratory of Urban ITS, Jiangsu Province Collaborative Innovation Center of Modern Urban Traffic Technologies, School of Transportation, Southeast University, city=Nanjing, postcode=210096, state=Jiangsu, country=China
1 Introduction
Graphs provide a flexible and expressive framework for modeling complex systems composed of interconnected entities. Formally, a graph $G=(V,E,X_{V},X_{E})$ consists of a set of nodes $V$, edges $E$, node attributes $X_{V}$, and edge attributes $X_{E}$, where both nodes and edges may carry heterogeneous features. Unlike grid-structured data such as images or sequences, graphs encode non-Euclidean topologies with irregular neighborhood structures and varying connectivity patterns. To support learning on such structures, graph learning aims to derive low-dimensional node or graph-level representations $Z\in\mathbb{R}^{d}$ that preserve both structural dependencies and semantic information from the original graph [1, 2, 3]. Due to these representational advantages, graph learning has been successfully applied across a wide range of domains, including social network analysis [4], personalized recommendation [5], transportation optimization [6], financial modeling [7], and bioinformatics [8]. As a result, it has become a foundational technique for extracting knowledge from relational data and enabling models to reason over structured domains.
Despite this progress, real-world graph learning faces significant challenges that diverge from idealized assumptions often made in academic benchmarks. For instance, in social networks, cold-start users may lack sufficient links or profile features, making it difficult to learn reliable representations [5]. In fraud detection, the data is often highly imbalanced, where fraudulent behavior accounts for only a small fraction of all activities, and labels are sparse and noisy [9]. In smart cities, traffic systems involve heterogeneous and dynamically evolving road networks, posing difficulties for models to adapt to changing topologies in real time [6, 10, 11]. As summarized in Figure 1, these concrete scenarios highlight four fundamental challenges in real-world graph learning:
Figure 1: The four fundamental challenges emerge of real-world graph complexity: (1) Incompleteness in graphs, where nodes, edges, or attributes are missing, (2) Imbalance in graphs, where the distribution of nodes, edges, or labels is highly skewed, (3) Cross-domain heterogeneity in graphs, where graph data from different domains exhibit semantic and structural discrepancies, and (4) Dynamic instability in graphs, where graphs undergo dynamic changes in topology, attributes, or interactions over time.
(1) Incompleteness in Graphs An incomplete graph refers to a graph where some node or edge information is missing, resulting in an incomplete representation of the structure or attributes that fails to fully reflect the complexity of the real-world system [12]. This challenge typically arises from incomplete data collection, missing link information, or outdated knowledge. For example, in social networks, missing user attributes—such as age, location, or interests—can obscure community boundaries, making it difficult to accurately detect user groups or recommend content. Similarly, missing connections between users may conceal important social ties, weakening the structural signals used for clustering or influence analysis. This incompleteness undermines the effectiveness of graph learning.
(2) Imbalance in Graphs A graph is considered imbalanced when certain categories or subgraphs contain significantly more nodes or edges than others [13]. This can lead to some categories dominating the graph, while others are left underrepresented [14]. For example, financial transaction networks are characterized by containing a large number of legitimate transactions, and only a small fraction of fraudulent ones [15]. Imbalanced training data may cause models to overfit to the majority class, thus impacting the model’s ability to generalize and recognize the minority class effectively [16]. This data imbalance challenge affects many graph learning methods, especially those relying on graph convolutional networks [17].
(3) Cross-Domain Heterogeneity in Graphs Cross-domain heterogeneity refers to significant disparities in attributes and structural patterns that arise when graph data are constructed from multiple source domains, especially when these domains have different data modalities or exhibit substantial distribution shifts. For instance, road network data collected from various cities may exhibit considerable heterogeneity, with some cities having grid-like structural patterns, while others exhibit radial configurations [18]. This issue becomes even more critical in the rapidly growing field of graph foundation models [19], which aim to develop generalizable models capable of handling diverse graphs from different domains, such as molecular graphs and citation networks. Cross-domain heterogeneity poses additional challenges for graph learning, as diverse attributes and structural patterns cannot be easily processed in a unified manner [20, 21]. Additionally, domain discrepancies may distort essential semantic and structural signals, complicating the identification of transferable features and limiting the effectiveness of graph learning methods [22, 23].
(4) Dynamic Instability in Graphs Dynamic instability in graphs refers to continuous and unpredictable changes in node features and relationships over time. These may include updates to node attributes, alterations in graph structure, or the emergence of new node and edge types. Structural changes often involve adding or removing connections or reorganizing how nodes interact. For example, in a large-scale knowledge graph like Wikidata, an entity initially labeled as an “athlete” might later become classified as a “coach”. This transition entails significant shifts in both node features and their connections within the graph [24, 25]. Furthermore, these continuous transformations introduce instability, making it difficult for graph learning models to adapt effectively. As a result, performance may degrade in tasks such as node classification, link prediction, and node representation learning [26]. A deeper understanding of these forms of dynamic instability can facilitate the development of more robust and adaptive graph learning methods, crucial for accurately modeling the evolving real-world graphs.
A variety of traditional graph learning methods have been developed to address these challenges, achieving notable success in specific contexts. For instance, graph completion algorithms [27] and graph autoencoders [12] have been employed to handle missing data; techniques like GraphSMOTE [28] and re-weighted loss functions [29] are commonly used to mitigate data imbalance; and graph domain adaptation methods, including adversarial regularization [20, 30] and distribution alignment [23, 31], have been explored to address cross-domain heterogeneity. Additionally, dynamic graph models such as EvolveGCN [32] and DyRep [33] have been proposed to model evolving graph structures. While these methods are effective within their respective scopes, they often rely on task-specific designs and require significant domain expertise.
In contrast, the emergence of large language models (LLMs) has opened up new opportunities for addressing the fundamental challenges of graph learning in a more unified and flexible way [34, 35, 36]. With their strong representational capacity, contextual reasoning abilities, and generalization potential, LLMs can extract rich semantic patterns from heterogeneous and noisy data, offering a promising complement—or even alternative—to traditional graph-specific approaches. To reflect this growing interest, Figure 2 shows publication trends over the past decade (up to May 2025) related to “Graph Learning” and “LLM-based Graph Learning”. The statistics indicate that not only has graph learning become an increasingly important topic, but the proportion of research exploring the integration of LLMs into graph learning has also been steadily rising.
Figure 2: Annual publication trends in Graph Learning from January 2015 to May 2025, with a highlighted subset of works related to LLMs. The blue bars represent the total number of publications in Graph Learning each year, while the overlaid pink segments indicate the number of publications that are specifically related to LLMs. The percentages above each bar denote the proportion of LLM-related papers within the total Graph Learning publications for that year. Data sourced from Google Scholar (https://scholar.google.com).
This survey provides a comprehensive review of current research at the intersection of graph learning and large language models, centered on the four fundamental challenges introduced above. To ensure broad coverage, we performed an extensive literature search across major databases (arXiv, ACM Digital Library, IEEE Xplore, Elsevier, and Springer, etc.) using keyword combinations such as “LLMs”, “graph learning”, “incompleteness”, “imbalance”, “heterogeneity”, “dynamic graphs”. We focus on recent publications (primarily 2015–2025) in top machine learning, data mining, and NLP venues, as well as relevant preprints. This survey includes papers that meet at least one of the following criteria: (1) the work explicitly addresses one of the four fundamental challenges in graph learning, or (2) the proposed solution incorporates LLMs as a central component. Based on these criteria, we initially collected and reviewed over 1,000 relevant papers. We then categorized them according to the type of challenge addressed and the strategies employed, ultimately selecting a curated set of over 260 representative works. For each challenge, we briefly discuss conventional (non-LLM) techniques to provide background context, while our primary focus is to highlight the novel contributions and advantages introduced by LLM-based approaches.
The structure of the paper is organized as follows: Section 2 provides an analysis and summary of traditional methods and LLM-based approaches for handling graph learning challenges. Section 3, building on the existing body of research, proposes new technical paths and solutions. Section 4 discusses the potential future directions for graph learning research. Finally, Section 5 concludes the paper.
2 Overview of Graph Learning and LLMs
Figure 3: Comparison between traditional graph representation learning and emerging paradigms that leverage LLMs. (a) Conventional graph learning relies on GNNs to encode structural information into node or graph-level representations. (b) In LLM-based approaches, LLMs can enhance graph structures using auxiliary text (top), extract semantic representations directly from graph inputs (middle), or generate task-specific answers through reasoning over graph structures (bottom).
2.1 Graph Learning
Graph learning refers to a family of machine learning techniques designed specifically to analyze and learn from graph-structured data [1, 37, 2]. Unlike data in Euclidean domains, graphs encode relationships between entities, and thus graph learning algorithms must leverage both node features and the graph topology. The goal is often to learn vector representations for nodes, edges, or entire subgraphs/graphs that capture their structural role and feature information, so that these embeddings can be used for downstream tasks, as shown in Figure 3(a).
Random-Walk-Based Embedding Methods: Early approaches to graph learning draw from analogies to signal processing and dimensionality reduction. For example, graph signal processing (GSP) methods [38, 39] treat node features as signals on a graph and use spectral decomposition of the graph Laplacian to filter or extract features. Techniques like graph Fourier transforms and wavelets provide mathematical insight but often assume the graph is static and have high computational cost for large graphs. Another line of work uses matrix factorization, such as factorizing adjacency or Laplacian matrices [40] or methods like LINE [41] that factorize higher-order proximity matrices, to embed nodes in a low-dimensional space. These capture global graph structure but might miss local detail and again can be limited by data sparsity or missing entries.
A breakthrough in graph learning came with the introduction of random-walk and neural embedding methods. Algorithms like DeepWalk [42] and Node2Vec [43] perform truncated random walks on the graph to generate sequences of nodes, then apply techniques from word2vec to learn node embeddings where co-occurring nodes in walks have similar vectors. These methods, along with variants like metapath2vec [44] for heterogeneous graphs, are scalable and capture community structure, but they treat the graph as a static corpus of walks and do not incorporate node attribute features directly.
Graph Neural Networks: The advent of graph neural networks (GNNs) marked a significant change by enabling end-to-end learning directly on graph structures. Spectral GNNs, starting with Bruna et al. [45], were initially limited by high computational costs, which were mitigated by subsequent works using polynomial approximations [46]. The introduction of Graph Convolutional Networks (GCNs) by Kipf and Welling [47] popularized neighborhood aggregation in semi-supervised settings. Further developments such as GraphSAGE [48], GAT [49], SGC [50], and GIN [51] improved the expressiveness, scalability, and adaptability of GNNs.
More recently, two powerful approaches have emerged in graph learning: contrastive learning and Transformer-based architectures. Graph contrastive learning methods like DGI [52], GraphCL [53], and HeCo [54] focus on learning robust representations by maximizing consistency across multiple views of the same graph. Transformer-inspired models, such as GT [55] and GraphGPS [56], introduce global self-attention and hybrid message-passing to model long-range dependencies and scale across graph types.
While graph learning techniques have shown remarkable success in modeling relational structures and topological dependencies, they often fall short in capturing high-level semantics and leveraging external unstructured data [57, 58]. These limitations become more pronounced when dealing with real-world graphs that are frequently incomplete, imbalanced, heterogeneous across domains, or dynamically evolving over time. These challenges have motivated a growing interest in integrating graph models with powerful language-based representations, such as LLMs, which offer strong semantic reasoning capabilities and the flexibility to incorporate external knowledge into the graph learning pipeline.
2.2 LLMs
LLMs are a class of deep neural networks pre-trained on massive corpora to model the probabilistic structure of natural language [59, 60]. By learning distributed semantic representations of tokens, LLMs are capable of capturing rich contextual dependencies, making them highly effective across a wide range of NLP tasks, such as text generation [61], question answering [62], machine translation [63], and sentiment analysis [64].
The evolution of LLMs represents a shift from manual feature engineering to end-to-end data-driven language understanding. Early NLP systems relied on symbolic rules and statistical models (e.g., regular expressions [65], HMMs [66], and CRFs [67]), which were limited by local context and domain specificity. The introduction of the Transformer architecture by Vaswani et al. [68] in 2017 enabled global attention-based modeling, laying the foundation for scalable deep language models.
Subsequent models such as BERT [69] introduced bidirectional context modeling through masked language modeling, while GPT-3 [70] demonstrated the power of the “Scaling Law” by achieving strong few-shot capabilities with 175 billion parameters. The introduction of ChatGPT [71], fine-tuned via reinforcement learning with human feedback (RLHF), brought conversational alignment and intent understanding to new levels of human-like interaction.
Recent models, including GPT-4 [72], Gemini [73], and DeepSeek-MoE [74], mark a new phase of innovation. These models integrate visual encoders, multimodal fusion, and mixture-of-experts (MoE) architectures, enabling cross-modal understanding while optimizing computation. DeepSeek-MoE, for instance, uses dynamic expert routing to reduce computational overhead without sacrificing accuracy, offering a scalable path to deployment.
LLMs, by virtue of their training, encapsulate a broad range of knowledge and have shown surprising abilities in reasoning, abstraction, and generalization. This makes them attractive for addressing graph challenges as follows:
Semantic reasoning: LLMs excel at inferring latent information and making logical connections based on context. They can understand descriptions of entities and predict relationships that are not explicitly stated but are likely given the context. For example, given a partial description of a node, an LLM might deduce other attributes or links that should exist.
Cross-modal and cross-domain integration: LLMs can serve as a bridge between different data modalities and domains. They can take textual input and integrate it with structural information by producing a unified representation or by generating structured outputs. This is highly useful for heterogeneous graphs – e.g., aligning a gene interaction network with a drug-side-effect network by reasoning over biomedical texts that link the two.
Generative augmentation: LLMs have the ability to generate content – whether it be new text, or by extension new graph elements described by that text. This generative power can create synthetic nodes or edges with plausible attributes for data augmentation, address class imbalance by generating additional samples, or even hypothesize new connections in knowledge graphs.
This background section has outlined the principles of graph learning and LLMs. Figure 3(b) illustrates representative paradigms for leveraging LLMs in graph learning, where LLMs interact with graph data in different roles to support structure enhancement, representation learning, and predictive reasoning. In the following section, we bring these together by examining how LLMs have been applied to each of the four fundamental graph learning challenges. We will see that for each challenge, LLMs introduce new ways to overcome the limitations of traditional graph algorithms, whether by injecting external knowledge, balancing data through generation, aligning heterogeneous information, or reasoning about changes.
3 Related Work and Current Techniques for Graph Learning Challenges
In this section, we discuss each of the four core challenges in detail. For each challenge, we first formalize the problem. We then review representative traditional and LLM-enhanced solutions to provide context.
3.1 Incompleteness in Graphs
In the real world, graph data is often incomplete due to missing node attributes, edges or labels (as shown in Figure 1), which significantly impacts the accuracy and generalization ability of graph models. By leveraging LLMs for graph data completion and the inference of missing information, the performance of graph learning models can be significantly improved, even in scenarios with incomplete or missing information. We categorize current approaches to incomplete graphs into three major directions (as shown in Figure 4): (i) Robust Graph Learning, which focuses on making models resilient to missing data; (ii) Few-Shot Graph Learning, which addresses scenarios with extremely limited labeled data or structure by transferring knowledge; and (iii) Knowledge Graph Completion, which is a prominent sub-area dealing with inferring missing facts in large knowledge graphs. The relevant references and categorization are presented in Table 1.
Figure 4: Illustrations of three representative tasks addressing graph incompleteness. Robust Graph Learning (left) handles noisy or missing attributes to ensure stable performance. Few-shot Graph Learning (middle) aims to generalize from limited labeled nodes or subgraphs. Knowledge Graph Completion (right) focuses on inferring missing links between entities based on observed relationships.
Table 1: Literature Summary of LLM-based Methods for Handling Incompleteness in Graphs.
Domain Incomplete Type Reference (Method—Citation) Robust Graph Learning Node LLM4NG[75], LLM-TAG[76], SPLLM[77] Label LLMGNN[78] Mixed GraphLLM[79], PROLINK[80], UnIMP[81] Few-Shot Graph Learning Structure LinkGPT[82], AnomalyLLM[83] Mixed LLMDGCN[84], HGT[85] FlexKBQA[86], KGQG[87] Knowledge Graph Completion Node LLM-KGC[88], GS-KGC[89], GLTW[90] Label KGs-LLM[91] Mixed FSKG[92], KGLLM[93], KICGPT[94] RL-LLM[95], GoG[96], KoPA[97] LLMKG[98], DIFT[99] CP-KGC[100], MuKDC[101]
3.1.1 Robust Graph Learning
Robust graph learning methods aim to maintain model performance even when the input graph is noisy or partially observed. Traditional GNNs often degrade in accuracy if key attributes are missing or if the graph is sparsely connected, because the iterative message passing has less information to propagate. Therefore, researchers have proposed specialized techniques to make graph learning robust to incompleteness.
(1)Traditional Methods for Robust Graph Learning Early efforts in robust graph learning focused on imputing missing values using global interpolation techniques under low-rank assumptions. For example, matrix completion methods [102, 27, 103, 104] estimate missing attributes based on the global structure of the data. GraphRNA [103] introduces Attribute Random Walk (AttriWalk), modeling attributes as a bipartite graph to capture attribute-structure interactions, thereby enhancing local adaptability and robustness in heterogeneous environments. ARWMF [104] combines random walks with matrix factorization and mutual information to improve unsupervised node embedding under noisy conditions.
With the rise of deep learning, more sophisticated generative models were developed to overcome the linear assumptions of early methods [12, 105, 106, 107, 108, 109, 110, 111, 112, 113]. SAT [12] enhances robustness by decoupling structure and attribute signals via shared latent spaces and distribution alignment. SVGA [105] imposes strong probabilistic regularization on latent variables to prevent overfitting when features are sparse or unreliable. Models like ITR [106] and VISL [107] progressively refine latent variables or inter-variable structures using topological cues, improving resilience against structural noise. More recently, methods like MATE [108] and AIAE [109] exploit multi-view and multi-scale generation strategies to stabilize representation learning under incomplete or corrupted input. CSAT [110] incorporates contrastive learning and Transformers to detect communities under noisy or weak supervision, further enhancing robustness through auxiliary signals.
Another line of work improves model robustness through dynamic inference mechanisms that rely on structural priors like homophily and community consistency [114, 115, 116]. For instance, FP [114] integrates Dirichlet energy minimization with graph diffusion to achieve stable feature recovery. PCFI [115] assigns confidence scores to feature channels, allowing uncertainty-aware propagation. ARB [116] addresses the cold-start problem by introducing virtual edges and redefined boundary conditions to improve propagation in sparse or poorly connected graphs. A notable advantage of such propagation-based methods is their parameter-free nature, making them highly adaptable and easily integrated with LLM-based pipelines for downstream reasoning tasks. To tackle the cold-start problem in node representation learning, Cold Brew [117] introduced a feature contribution ratio metric to guide teacher-student distillation, uniquely encoding topological sparsity as a dynamic temperature coefficient. CTAug [118] employs subgraph-aware contrastive learning to preserve dense subgraph priors, significantly enhancing representation learning in dense graphs.
These methods significantly enhance the predictive capabilities of graph data in scenarios with missing attributes by jointly modeling, employing variational inference, and optimizing multi-source information. Nevertheless, three challenges remain: (1) coupling optimization of completion and prediction dramatically increases training complexity; (2) latent-space completion methods are sensitive to prior assumptions; (3) efficient incremental mechanisms are lacking for dynamic attribute completion.
(2) LLM-enhanced methods Recently, researchers have begun exploring the deep integration of semantic reasoning and incomplete graph learning. Methods based on LLMs [76, 75, 77, 79, 80, 78] have pioneered a new approach of “semantic completion–topology optimization” co-evolution, leveraging the strong generative and reasoning capabilities of LLMs.
For instance, Sun et al. [76] proposes an LLM-based architecture that effectively tackles graph incompleteness by leveraging semantic reasoning to infer missing edges and nodes. Specifically, LLMs can identify and establish plausible connections between nodes based on their semantic similarity, thereby mitigating the common challenge of missing or incomplete relationships in real-world graphs. Experimental results indicate that these LLM-enhanced methods surpass traditional approaches—such as basic link prediction or imputation techniques—by producing more accurate and semantically meaningful graph structures. Similarly, LLM4NG [75] employs LLMs to generate contextually relevant nodes, enriching sparse or incomplete graph data, particularly in few-shot learning scenarios. This approach significantly improves graph completeness by addressing gaps that conventional methods (e.g., imputation) often fail to resolve effectively.
Meanwhile, SPLLM [77] tackles the challenge of incompleteness in high-dimensional sensor graphs, where missing values disrupt both spatial and temporal dependencies. It integrates a spatiotemporal fusion GCN with LLM-based fine-tuning, using external domain knowledge (e.g., road networks, weather data) to infer missing patterns and complete partial features more effectively. UnIMP [81], in a similar spirit, targets mixed-type tabular data by leveraging LLMs with bidirectional high-order message passing on cell-oriented hypergraphs. This approach addresses incompleteness arising from semantic and structural heterogeneity, which traditional imputation methods struggle with. Figure 5 illustrates how LLMs contribute to graph completion by extracting domain-specific semantic features from textual descriptions or external knowledge. These features serve as compensatory signals for missing nodes, edges, or attributes in sparse or incomplete graphs.
To further close the semantic gap introduced by graph incompleteness, several works propose end-to-end frameworks combining LLMs and graph models. For instance, GraphLLM [79] introduces a unified architecture where a graph encoder maps node features into the LLM’s semantic space, enabling it to reason over incomplete or noisy node attributes using attention-gated fusion of structure and context. This helps, for example, in citation networks or recommendation systems where node metadata may be sparse or outdated. PROLINK [80] enhances inductive reasoning on low-resource knowledge graphs by generating structural prompts and using LLMs to fill in gaps from limited graph and text data — particularly valuable in biomedical or emerging domains with sparse curation. Lastly, LLM-GNN [78] proposes a semi-supervised approach where LLMs annotate a small subset of nodes, helping GNNs generalize from minimal labeled data — addressing the classic label incompleteness challenge in large-scale graphs.
Empirical results in the literature indicate that LLM-augmented models achieve higher accuracy on incomplete graph benchmarks, especially when missingness is significant.
Figure 5: The framework addresses incomplete graph structures by integrating structural signals from GNNs and semantic priors from LLMs. The “Align / Fusion / Distillation” module enables the model to compensate for missing graph information by leveraging textual context, resulting in enriched and unified node attribute/edge/label representations
3.1.2 Few-shot Graph Learning
Few-shot learning (FSL) deals with scenarios where very few labeled examples are available for training, or where a graph has only a handful of nodes/edges in a particular class or subset of interest. This is related to incompleteness in that labels can be viewed as a type of missing information.
(1) Traditional Methods for Few-Shot Learning FSL aims to enable effective model generalization from a limited number of labeled samples. In graph domains, FSL not only addresses label scarcity but also tackles structural challenges such as missing links and infrequent or long-tail relations [119, 120, 121, 122, 123]. To overcome these limitations, traditional approaches focus on explicitly modeling graph structures and relational semantics under low-resource conditions.
For example, Gmatching [119] pioneered the use of single-hop neighborhood structures and relation prototype matching to handle rare relation prediction. GFL [120] establishes a transferable metric space by projecting auxiliary knowledge onto the target graph. FSRL [121] proposes a heterogeneous relation meta-learning framework with a reference representation generator to decouple relation semantics. FAAN [122] introduces an adaptive encoder and aggregator to model entity-reference contributions, significantly enhancing few-shot knowledge graph completion through fine-grained semantic representations. For real-world applications, MetaHG [123] addresses drug trafficking detection on Instagram by constructing a heterogeneous graph and optimizing node representations via meta-learning, outperforming existing methods.
Despite these advancements, existing methods often rely on hand-crafted architecture designs and still face inefficiencies in handling complex graph structures, along with limited capabilities for structured and interpretable reasoning.
(2) Few-Shot Learning in LLMs Few-shot learning within LLMs has seen remarkable progress, particularly in tasks such as graph data analysis [84], link prediction [82], anomaly detection [83], table understanding [85], and knowledge graph question answering (KGQA) [86, 87].
LLMDGCN [84] introduces a degree-aware prompt tuning framework, encoding graph structures as positional embeddings for LLMs. This method not only filled missing node labels but also restored missing edges through an edge prediction module, improving classification accuracy and robustness in low-label scenarios. For link prediction in graph data, LinkGPT [82] extends LLMs’ few-shot learning capabilities by incorporating instruction tuning and retrieval-based re-ranking mechanisms. This approach significantly improved model reasoning under limited training data, yielding substantial performance gains on real-world datasets. In anomaly detection for dynamic graphs, AnomalyLLM [83] combines LLM pretraining knowledge with few-shot learning strategies, using a dynamic-aware encoder to generate edge representations and incorporating word embedding prototypes for edge reprogramming, successfully enhancing few-shot anomaly detection effectiveness. For table understanding tasks, the HGT framework [85] utilizes heterogeneous graph-enhanced LLMs, integrating soft prompts and instruction tuning to effectively align table semantics with LLM parameter knowledge, leading to more efficient table parsing in few-shot scenarios. In KGQA, FlexKBQA [86] improves question-answering performance in low-resource settings by converting knowledge base programs into natural language questions, enabling synthetic data generation. Meanwhile, the zero-shot KGQA framework [87] combines LLMs with smaller models, significantly reducing generation costs while achieving performance comparable to fully supervised models.
In summary, while few-shot graph learning remains challenging, LLMs act as few-shot learners by nature – they can leverage their pre-trained knowledge to make sense of novel tasks with limited data. Results show that even with minimal graph information, an LLM can propose connections or classifications that align with human knowledge, essentially providing a powerful prior to the graph model.
3.1.3 Knowledge Graph Completion
Incompleteness is inherent in knowledge graphs: no knowledge base is ever complete. The development of knowledge graph completion (KGC) revolves around the core challenge of structure-semantics fusion, forming a progressive innovation trajectory from fundamental approach exploration to multimodal collaboration. LLMs, with their strong language understanding and reasoning capabilities, demonstrate great potential in addressing long-tail challenges, reducing annotation burdens, and handling incomplete graph data [91, 92, 93, 95, 96, 88, 97, 101, 98, 99, 100, 89, 94].
Early studies focused on converting knowledge graph triples into textual sequences to validate the feasibility of LLM-based generation for KGC. FSKG [92] introduces an LLM-based generation and refinement framework to address data scarcity in long-tail scenarios. By employing multi-stage knowledge distillation, MuKDC effectively generates supplementary knowledge, improving KGC performance. KGs-LLM [91] proposes an innovative knowledge graph generation method utilizing generative LLMs (e.g., GPT-3.5). By leveraging zero-shot prompting and a strategy that requires no external knowledge, this approach iteratively extracts graph components without manual annotation, making it highly scalable and effective in domain-specific datasets. Similarly, KGLLM [93] transforms knowledge graph triples into text sequences and leverages LLMs for prediction, achieving state-of-the-art performance in triple classification and relation prediction. Compared to traditional methods, this approach enables efficient fine-tuning of smaller models, particularly lightweight LLMs such as LLaMA-7B. To mitigate the long-tail entity issue in KGC, KICGPT [94] integrates an LLM with a triple-based KGC retriever, encoding structural information into prompts using contextual learning strategies. This method requires no additional training costs and performs exceptionally well in few-shot settings.
To address the structural information loss inherent in text-based methods, researchers have developed multi-level prompt encoding techniques. RL-LLM [95] applies few-shot learning and multi-prompt optimization for relation prediction in e-commerce knowledge graphs, achieving performance comparable to human-labeled data in some cases, showcasing the potential of LLMs in such tasks. GoG [96] introduces a ”think-search-generate” approach, enabling the generation of new triples without additional training, thereby improving incomplete knowledge graph reasoning. This framework effectively integrates LLMs with knowledge graphs, significantly enhancing question-answering over incomplete KGs. LLM-KGC [88] embeds ontology knowledge and graph structure into LLM prompts, combining topological information with contextual learning techniques to achieve effective KGC across different settings. Experiments show that this approach outperforms baseline models across multiple datasets, particularly excelling in inductive reasoning tasks where strong inference capabilities are required. KoPA [97] introduces a knowledge prefix adapter, incorporating structural information into the LLM inference process. By converting graph entities and relations into textual representations suitable for LLM input, KoPA significantly enhances KGC performance, strengthening structure-aware reasoning in KGC tasks.
Building upon the advancements in LLM-driven KGC, several frameworks have further optimized few-shot KGC (FKGC), entity alignment, prompt engineering, and structural integration to enhance performance and scalability. MuKDC [101] employs multi-level knowledge distillation to generate supplementary knowledge while maintaining coherence and reliability, excelling in long-tail FKGC tasks through multimodal knowledge generation and consistency evaluation. LLMKG [98] enhances zero-shot and one-shot KGC using Mixtral-8x7b and GPT-4, demonstrating that well-designed prompts allow LLMs to perform effectively under various evaluation metrics. DIFT [99] improves entity alignment by integrating lightweight models and discriminative prompts, preventing alignment errors while significantly boosting KGC results. CP-KGC [100] enhances semantic richness in quantized LLMs through contextually adaptive prompts, effectively handling polysemous entities and pushing the boundaries of text-based KGC. Beyond textual enhancements, structural integration methods such as GLTW [90] and GS-KGC [89] further refine KGC. GLTW improves local and global structure encoding by integrating Graph Transformers with LLMs, while GS-KGC enhances contextual reasoning by leveraging subgraph-based question-answering and negative sampling, significantly improving LLM accuracy in missing triple identification.
Overall, LLMs bring in a semantic understanding of the KG elements, which is valuable because KGs often have textual labels for entities/relations that carry meaning. LLMs leverage this to make more informed predictions, especially when data is sparse or the pattern is not purely structural but semantic.
3.1.4 Summary of Incompleteness
Across robust learning, few-shot learning, and KGC, the common thread is that LLMs serve as knowledge-infusers and intelligent guessers for what is missing in the graph. They help create a more complete picture by either filling in data directly or guiding the graph model on where to look. Empirically, studies report that incorporating LLM-generated features or suggestions leads to substantial gains in tasks like node classification with missing features, link prediction with sparse edges, and KG completion benchmarks. As LLMs continue to improve, especially in domain-specific knowledge, we expect their role in handling graph incompleteness to grow.
3.2 Imbalance in Graphs
Graph data often exhibits imbalance that can severely affect the performance, particularly in tasks such as node classification (as shown in Figure 6). This imbalance typically complicates the learning process as models tend to be biased toward the majority class or low-degree nodes, resulting in poor generalization for underrepresented elements. To address this, techniques such as resampling, re-weighting, and graph-based regularization have been proposed to alleviate the effects of data imbalance. However, these approaches often require domain-specific adjustments and may not always yield optimal results.
Figure 6: Illustration of three representative tasks addressing graph imbalance. Class-imbalanced graph learning (left) refers to scenarios where certain node classes (minority classes) are underrepresented compared to others. Structure-imbalanced graph learning (right) highlights disparities in node connectivity, where high-degree nodes dominate information flow while low-degree nodes are underexposed.
Recent advances in LLMs offer promising solutions to the challenges posed by imbalanced graph data. LLMs, with their ability to process and generate semantically rich representations, can be leveraged to enrich the graph’s feature space, enabling more nuanced and balanced learning.
This section reviews the application of traditional methods and LLMs for learning from imbalanced graph data. The literature is categorized into two key research areas: Class Imbalance and Structure Imbalance. The relevant references and categorization are presented in Table 2.
Table 2: Literature Summary of LLM-based Methods for Handling Imbalance in Graphs.
Domain Task Reference (Method—Citation) Class Imbalance Graph Learning Node Classification LLM4NG[124], LLM-GNN[78], G2P2[125] LA-TAG[126], GSS-Net[127], TAGrader[128] LLM-SBCL[129], SEGA[130], SocioHyperNet[131] Cella[132], LLM-TIKG[133], ANLM-assInNNER[134] LLM-HetGDT[135] Prediction LKPNR[136], LLM-DDA[137] Graph Completion KICGPT[94], KGCD[138] Foundation Model GraphCLIP[139] Structure Imbalance Graph Learning Node Classification GraphEdit[140] Graph Completion SATKGC[141], MPIKGC[142], LLM4RGNN[143]
3.2.1 Class-Imbalanced Graph Learning
Class imbalance refers to scenarios where the sample size of certain categories is significantly smaller than others. For instance, in social networks, ordinary users vastly outnumber influential “key opinion leaders”, while in fraud detection systems, legitimate transactions overwhelmingly dominate over fraudulent ones. Traditional classifiers, including GNNs, often exhibit a bias toward predicting majority-class samples, resulting in suboptimal recognition accuracy for minority classes [16, 13]. Furthermore, GNNs update node representations by aggregating neighborhood information. Minority-class nodes risk being structurally homogenized or misrepresented when their local neighborhoods are dominated by majority-class nodes, leading to a propagation of representation bias across the graph.
(1) Traditional Methods Traditional methods for addressing class imbalance in graph data primarily include clustering, contrastive learning, data augmentation, and innovative loss function. Most articles combine one or more of the above techniques to solve the class imbalance problem.
ECGN [144] proposes a cluster-aware graph neural network framework that explicitly models the graph cluster structure to balance class representations, utilizing intra-cluster contrastive loss to enhance the distinguishability of minority class nodes. Similarly, C3GNN [145] clusters majority class nodes into subclasses and contrasts them with the minority class, using Mixup techniques to enhance the semantic substructure representation and alleviate the class imbalance problem. These methods share a key insight: clustering helps balance the class distribution by dividing majority class samples into multiple subclasses, and contrastive learning can improve the separability of minority nodes by constructing positive and negative sample pairs. Building on this, CCGC [146] designs a clustering-based positive sample selection strategy and optimizes graph representations through hierarchical contrastive learning, reducing overfitting to the minority class. ImGCL [147] introduces a progressive balanced sampling strategy, dynamically adjusting contrastive weights based on node centrality measures, and utilizes pseudo-labeling to address topological imbalance. These works highlight how clustering and contrastive learning can be synergistically combined to tackle both structural and label imbalance.
Another line of work focuses on data augmentation, which involves generating diverse topologies or attribute changes for minority classes to balance the data distribution. GraphSMOTE [28], a pioneering method, interpolates minority nodes in the embedding space and predicts edges to preserve topological consistency. However, such interpolation-based methods may struggle with semantic coherence. To address this, GraphSHA [148] generates hard samples via feature mixing and employs a semi-mixed connection strategy to prevent neighbor-class encroachment, enhancing boundary discrimination. Augmentation is often paired with adaptive loss functions. For example, TAM [149] introduces topology-aware margin loss, dynamically adjusting classification boundaries based on local neighbor distributions of nodes, addressing decision bias caused by label node location shifts.
In addition, HGIF [150] addresses the heterogeneous graph scenario by using invariance learning to separate semantically invariant features, enhancing the robustness of fraud detection in out-of-distribution settings. GraphSANN [151] breaks the homophily assumption and designs a topological reconstruction framework to tackle the class imbalance problem in heterogeneous graphs.
(2) Methods Based on LLMs LLM-based techniques offer a novel approach to address class imbalance. Researchers can overcome the limitations of traditional methods by leveraging three aspects: semantic-driven data augmentation, context-aware text understanding, and external knowledge injection. Figure 7 illustrates how LLMs can address class imbalance in graph learning by synthesizing semantically rich representations for minority class nodes. In this framework, graph neighborhoods, node attributes, and type constraints are linearized into natural language prompts, serving as the sole input modality for the LLM. This allows the model to generate new nodes or edge candidates with both semantic coherence and structural compatibility.
Figure 7: Illustration of an LLM-based pipeline for addressing class imbalance in graphs. Minority class information is first extracted and linearized into textual prompts, which are processed by the LLM to generate semantically enriched nodes and edges. The newly synthesized elements (highlighted in red) are integrated into the original graph to enhance representation and balance class distributions.
First, some methods utilize the text generation capabilities of LLMs for node attribute descriptions (such as entity descriptions and label semantic expansion) to synthesize more semantically coherent features for minority class nodes. LA-TAG [126] proposes a framework for text attribute rewriting and synthesis based on LLMs. It generates minority class node descriptions compatible with graph topology through instruction tuning, solving the semantic disconnection problem in text-graph alignment of traditional methods. Similarly, LLM4NG [124] designs a node generation approach for few-shot scenarios. It utilizes the LLM decoder to generate virtual nodes with both semantic coherence and structural consistency, while applying contrastive regularization to constrain the embedding space distribution. The method in [127] also integrates graph structure understanding modules with LLM generators to achieve fine-grained synthesis of sparse data streams, dynamically balancing class distributions in real-time graph data.
Another approach focuses on the contextual understanding and reasoning capabilities of LLMs, mining implicit relationships of minority classes through structure-aware reasoning. Both LLM-GNN [78] and Cella [132] proposed a zero-shot classification framework based on LLM semantic reasoning. By parsing node attributes and neighbor context, they directly generate category prediction confidence and get rid of the dependence on labeled data. G2P2 [125] introduces a graph-guided pre-training prompt framework, using LLMs to generate augmented text related to graph structure, enhancing the discriminability of low-resource classes in heterogeneous spaces. Meanwhile, LLM-TIKG [133] builds a threat intelligence knowledge graph, using LLMs to extract attack pattern association rules from unstructured text, and enhancing the representation of minority class threat entities through heterogeneous graph attention.
Other methods focus on external knowledge injection and transfer, mining long-tail information from open domains or domain-specific knowledge to improve the generalization ability of the model. KICGPT [94] dynamically builds prompt templates aligned with the graph structure based on contextual knowledge injection, enhances the reasoning ability of long-tail relations, and significantly improves the knowledge graph completion performance of sparse categories. LKPNR [136] combines knowledge graphs with LLM user preference reasoning to solve the cold start problem in long-tail news recommendation through semantic path enhancement. KGCD [138] uses LLM to generate pseudo-triplets based on logical rules to guide the knowledge graph completion model to focus on long-tail relations. At the same time, the method in [128] proposes a hierarchical knowledge distillation framework to transfer the open-domain semantic knowledge of LLM to the graph encoder to improve the generalization ability in the scenario of uneven categories. GraphCLIP [139] uses the cross-modal mapping capability of LLM to construct a graph-text comparison alignment framework to alleviate the problem of uneven categories in the target domain. Finally, LLM-DDA [137] integrates the disease-drug mechanism representation obtained by LLM reasoning into graph reconstruction in the biomedical scenario, and improves the accuracy of drug repurposing prediction for rare diseases through a multimodal attention mechanism.
In addition to the aforementioned new technologies, the introduction of LLMs has accelerated innovation in cross-domain applications. SEGA [130] uses LLMs to parse the structured elements of clinical conversation graphs, extracting depression semantic clues and social interaction patterns, enhancing the recognition of minority positive samples in mental health classification. The method in [131] proposes an LLM-enhanced sociological analysis framework based on hypergraphs, addressing behavior class imbalance in social media data through personality trait hyperedge modeling and LLM semantic deconstruction. ANLM-assInNNER [134] develops an automatic construction system for robot fault diagnosis knowledge graphs, using LLMs to generate fine-grained fault entity descriptions and balancing the data distribution of equipment status categories. LLM-SBCL [129] combines graph neural networks with LLM cognitive state modeling, improving the reliability of predicting questions on less common knowledge points through semantic enhancement of learner-question interaction graphs.
In summary, LLM-based approaches for class imbalance aim to even out the information content per class by generating additional data or highlighting distinguishing features, rather than just mathematically re-weighting or duplicating existing data.
3.2.2 Structure-Imbalanced Graph Learning
Structural imbalance refers to skewness in graph topology that can hinder learning. A typical example is a hub node vs. peripheral node issue: Hub nodes with very high degree can dominate aggregation and also often have many more training signals, whereas low-degree nodes might be ignored or get a very noisy aggregate from a single neighbor. Similarly, certain substructures might be over-represented. For example, in a molecule graph dataset for drug discovery, maybe most molecules contain a benzene ring but only a few contain a rarer motif; a GNN might mostly learn features relevant to benzene rings and be less sensitive to the rare motif, which could be critical for certain properties.
(1) Traditional Methods The main techniques for addressing structural imbalance in graph data are structural enhancement or debiasing. Some methods are dedicated to dynamically adjusting the neighbor aggregation weights to weaken the over-dominance effect of high-centrality nodes. DegFairGNN [152] introduces a generalized degree fairness constraint that reallocates neighborhood attention weights between high- and low-degree nodes to correct structural biases in GNN message passing, alleviating the issue of central nodes overly dominating the process. In addition, KG-Mixup [153] analyzes degree bias in knowledge graph completion, proposing a degree-aware contrastive loss function to dynamically adjust geometric constraints in the entity embedding space for high- and low-degree entities. Another approach starts from the perspective of subgraph structure enhancement and improves the expressiveness of the tail structure by generating diversified substructures. SOLT-GNN [154] designs a size-aware hierarchical pooling strategy that balances the representation distribution of graph structures of different sizes through subgraph cropping and feature decoupling, addressing the long-tailed structural imbalance in graph classification. SAILOR [155] introduces a structural enhancement-based framework for tail node representation learning, enhancing the visibility of low-degree nodes in the global topology through k-hop subgraph expansion and adversarial edge generation strategies. There are also methods that alleviate the number and topological imbalance of graph structures from the perspective of global structure reconstruction and regularity constraints. HiRe [156] constructs a hierarchical relational meta-learning framework that employs a meta-path-guided negative sampling mechanism to balance the structural coverage density of head and tail relations in knowledge graphs. QTIAH-GNN [157] develops a heterogeneous graph neural network that is aware of both quantity and topological imbalance, addressing node number imbalance and topological connection pattern imbalance through meta-relation-specific neighbor sampling and topological entropy regularization.
Although existing methods have alleviated the problem of graph structural imbalance to a certain extent, they still have limitations. They often rely on manually designed constraint rules, which leads to limited generalization ability and difficulty in adapting to complex and changing graph structural deviations. At the same time, these methods tend to ignore the semantic associations of the global topology, and usually adopt static adjustment strategies, which cannot dynamically adapt to the structural evolution patterns implicit in the graph. These limitations [16, 13] highlight the need for a new technique that can take into account both semantic understanding and dynamic adaptability to more comprehensively solve the problem of graph structural imbalance.
(2) Methods Based on LLMs Compared to traditional methods, LLM-based methods can achieve semantically driven structure completion and enhancement by generating context-aware virtual edges or subgraph structures to supplement the topological connections of tail nodes, as shown in Figure 8. Moreover, LLMs deeply integrate semantic understanding and structural analysis, identifying the underlying semantic causes of structural imbalance through causal reasoning (e.g., the logical absence of long-tail relationships in knowledge graphs) and generating logically constrained suggestions, rather than relying solely on statistical distribution adjustments.
Figure 8: Structural Imbalance Remediation via LLM-Based Text-Topology Co-Optimization. The optimized graph (right) shows the rebalanced connectivity, with new connections indicated in red.
SATKGC [141] proposes a contrastive learning framework based on subgraph awareness, generating semantic embeddings aligned with local subgraph structures using LLMs. It also designs a structure-aware contrastive loss function to enhance the logical coherence of long-tail relationships in knowledge graphs. MPIKGC [142] integrates LLMs from three perspectives: semantic enhancement, structural correction, and logical reasoning. Through dynamic weight allocation and logical rule generation, it addresses the relationship distribution imbalance in knowledge graph completion tasks. GraphEdit [140] develops an LLM-based iterative graph optimization framework, utilizing a semantic generation-topology validation mechanism to dynamically correct edge connections (such as adding virtual edges for tail nodes or pruning noisy connections), achieving collaborative optimization of global structural balance and local semantic consistency. LLM4RGNN [143] distills the reasoning capabilities of GPT-4 into a local LLM to identify malicious edges and, in conjunction with a language-model-based edge predictor, recovers a robust graph structure.
LLMs can semantically enrich or complete graph structures, but they typically need the graph to be represented in a suitable textual form. The advantage is that they might introduce connections or structural patterns that are semantically plausible, not random, thereby improving minority structures in meaningful ways.
3.2.3 Summary of Imbalance
Traditional graph imbalance methods mitigate bias via re-sampling, cost adjustment, or graph augmentation, but they lack external semantic context. LLM-integrated methods aim to generate new graph content or features that specifically bolster the minority classes or structures. As reported in recent studies, using LLMs in this way can significantly improve classification performance on long-tail classes and ensure that even structurally unique nodes are recognized. Essentially, LLMs function as an intelligent oversampling mechanism: instead of naive duplication, they produce novel yet relevant samples in data space. The result is often more balanced training data for the graph model and better generalization to minority cases.
3.3 Cross-Domain Heterogeneity in Graphs
Real-world graph data is often collected from multiple source domains [158, 159, 160], which can exhibit significant heterogeneity, referring to extreme disparities in both attributes and structural patterns. This heterogeneity typically arises when the graph data is collected or integrated from domains with inconsistent data modalities or distributions. Cross-domain heterogeneity introduces significant challenges for graph analysis and modeling, as the severe disparities make it difficult to unify these data into a common representation space [20, 161] and even hinder the detection of valuable and transferable features that can generalize well throughout the graph [22, 23]. Moreover, addressing cross-domain heterogeneity is a key prerequisite for building graph foundation models for more generalizable and scalable graph learning across diverse real-world applications [162].
While traditional graph domain adaptation methods [163] and multimodal graph learning methods [164] can help bridge the gap between different distributions or modalities from different graph domains, they typically require large-scale training datasets from different domains and cannot generalize to some unseen domains during training. In contrast, LLMs, with their superior comprehension and generalization abilities, can extract valuable and transferable semantic and structural features within a unified representation space without training from scratch, effectively addressing the challenges posed by cross-domain heterogeneity in graph learning [165].
Figure 9: Illustration of three representative tasks addressing cross-domain heterogeneity in graphs. Text-attributed graph learning (left) involves graphs where nodes or edges are associated with textual content, requiring joint modeling of structure and language. Multimodal attributed graph learning (middle) integrates diverse data types across nodes or relations. Structural heterogeneous graph learning (right) with complex cross-modal dependencies and heterogeneous attribute spaces.
In this section, as illustrated in Figure 9, we categorize cross-domain heterogeneity in graph data into three types: within-modality attribute heterogeneity, cross-modality attribute heterogeneity, and structural heterogeneity. Since LLMs are specifically designed for modeling textual modality, our focus on within-modality attribute heterogeneity primarily centers around textual attributes. For each type, we first review traditional graph learning methods, analyzing their strengths and limitations, and then discuss recent research leveraging LLMs to unify graph data and address the corresponding challenges. The relevant references and categorization are summarized in Table 3.
Table 3: Literature Summary of LLM-based Methods for Handling Cross-Domain Heterogeneity in Graphs.
Type Domain Reference (Method—Citation) Text-Attributed Graph Learning Textual Attribute Alignment TAPE[166], LLMRec[167], MINGLE[168], GHGRL[169] Graph Foundation Model OFA[162], UniGraph[170], BooG[171], Hyper-FM[172] Multimodal Attributed Graph Learning MLLM-based Multimodal Alignment LLMRec[167], MAGB[21] Graph-Enhanced Multimodal Alignment MMGL[173], GraphAdapter[174], TouchUp-G[175] UniGraph2[176] Structural Heterogeneous Graph Learning Topological Graph Textualization LLMtoGraph[177], NLGraph[178], Talk-like-a-Graph[179] GPT4Graph[180], GITA[181], LLM4Hypergraph[182] Attributed Graph Textualization GraphText[183], WalkLM[184], Path-LLM[185] InstructGLM[186], MuseGraph[187], Graph-LLM[188] Graph Token Learning GNP[189], GraphToken[190], GraphGPT[191] LLaGA[192], TEA-GLM[193], HiGPT[194]
3.3.1 Text-Attributed Graph Learning
Textual attributes are common in real-world graph data, such as the abstract of each paper in a co-citation network [195] or item descriptions in a recommendation network [167]. Although these textual attributes appear in the same modality, they can demonstrate significant distribution heterogeneity due to different sources. For example, textual attributes such as clinical notes could come from different healthcare providers, each with their own writing style, vocabulary, and context [196]. Similarly, clinical notes might be informal and easy to understand for patients, while medical codes should be more formal and standardized. Moreover, the textual attributes can appear in different languages [197], further increasing the heterogeneity. The distribution heterogeneity in textual attributes can lead to inconsistent semantic representations, making it difficult for graph learning methods to effectively capture transferable features.
(1) Traditional Methods for Textual Attribute Modeling To handle the heterogeneity in textual data, early data integration methods extract structured information from unstructured textual inputs, thereby reducing variability across sources and producing unified features for downstream tasks [198, 199, 200]. However, these approaches require manually design of data schemes and training extraction models tailored to specific applications. In contrast, statistical methods, such as Bag of Words and TF-IDF [201], have been introduced to automatically generate unified feature vectors from text without relying on domain-specific design. These feature vectors are often used as inputs for graph learning methods, such as GNNs [47, 48, 49], which further incorporate structural information from the graph to obtain more effective node representations for addressing downstream tasks. For instance, TADW [202] approximates DeepWalk [42] using matrix factorization, where the TF-IDF feature vectors of node textual attributes serve as the initial feature matrix. Paper2vec [203] utilizes learnable Word2vec [204] text embeddings as initial node features, which are then trained by predicting whether two nodes belong to the same neighborhood in the graph. While these early statistical methods or shallow models offer solutions for mapping heterogeneous textual attributes into a unified space, their limited expressiveness cannot effectively capture complex features in these textural attributes. Moreover, these methods adopt text embeddings as fixed initial node features rather than integrating them with structural learning, limiting the potential of leveraging graph topology to enhance the textual attributes representation learning [195].
Several advanced methods explore approaches for modeling textual attributes that can better integrate with graph structures, aiming to improve the representation learning of both textual attributes and graph structure data simultaneously. GraphFormers [195] proposes a GNN-nested architecture where the transformer-based textual embedding modules and the GNN modules are nested and trained together. Heterformer [205] introduces virtual neighbor tokens that capture information from both text-rich and textless neighbors, allowing the model to simultaneously consider textual semantic information and structural information to embed the textual node attributes. GIANT [206] leverages a self-supervised learning task called neighborhood prediction to generate graph-dependent text embeddings, which are then used in downstream node classification tasks. E2EG [207] extends GIANT [206] by introducing an end-to-end model that simultaneously learns node classification and neighborhood prediction, improving the node classification performance and computational efficiency. METAG [208] employs a single pretrained language model to learn multiplex text representations, utilizing relation prior tokens to capture relation-specific signals while maintaining parameter efficiency.
Although these advanced methods can embed heterogeneous textual node attributes into a unified representation space, their dependence on complex deep architectures requires a sufficient amount of training data. In applications with limited data availability, the model may fail to capture the valuable features from excessive heterogeneous textual attributes, and the learned features might fail to generalize effectively to textual attributes from other sources.
(2) LLM-based Methods for Textual Attribute Modeling LLMs, as language models, are naturally suited for modeling attributes represented in textual modality. With advanced language comprehension and generalization capabilities, LLMs can effectively capture the semantic meanings from heterogeneous textual attributes across diverse source domains and project them into a unified representation space that preserves the semantic information.
Figure 10: Handling cross-domain textual attribute heterogeneity with LLMs. Attributes from different domains are first transformed into unified textual descriptions by an LLM, then converted into unified embeddings using a trainable small LM to generate initial features for graph learning methods.
A number of methods leverage the use of LLMs to generate unified textual descriptions, which can be fed into a trainable smaller language model (LM) to extract task-specific embeddings, as depicted in Figure 10. TAPE [166] leverages the powerful language comprehension capabilities of LLMs to predict the category of a node based on its textual attributes and generate explanations for the prediction, which can be regarded as enhanced and aligned textual outputs. These aligned textual outputs are fed into a smaller language model to generate node feature vectors. As shown in [165], the domain shift is reduced in the LLM-enhanced text compared to the original textual attributes. Similarly, LLMRec [167] leverages LLMs to summarize user profiles and item attributes within a recommendation network, reducing the heterogeneity of the original attributes and producing aligned textual representations. These representations are then reprocessed by LLMs to generate unified embedding vectors for both users and items. Instead of generating unified textual descriptions, MINGLE [168] utilizes LLMs to map clinical notes and medical codes into a unified embedding space and uses these unified embedding vectors as initial node features for training a hypergraph neural network. GHGRL [169] utilizes LLMs to unify textual attributes on heterogeneous graphs based on format types and content types. Specifically, it first summarizes types information of all nodes using LLMs, which are then used to generate attribute summaries for each node based on the predicted types. These generated attribute summaries are then fed into a language model to generate unified feature vectors for training a graph neural network.
By transforming categorical and numerical features into textual attributes with language-based descriptions, LLMs can extend their ability to handle the heterogeneity in these structured attributes, which do not naturally belong to the textual modality. For example, in molecular graph modeling, node attributes such as atom types or properties are typically represented as categorical or numerical values. By converting these attributes into descriptive text and even enriching with domain-specific explanations, LLMs can effectively understand their semantic meaning and unify them within a common representation space. Both OFA [162] and UniGraph [170] address attribute heterogeneity by constructing Text-Attributed Graphs (TAGs). These models enhance textual attributes by incorporating additional semantically rich contextual descriptions. Furthermore, they leverage domain knowledge to enrich non-textual attributes with textual representations, enabling LLMs to process diverse attributes in a unified manner and generate consistent embedding vectors. BooG [171] follows a similar approach but further employs a contrastive learning-based pre-training objective, which enhances the ability to learn expressive representations and generalize across different domains and downstream tasks. Similarly, Hyper-FM [172] leverages a language model to extract semantic features from cross-domain textual attributes on hypergraphs and integrates structural information through hierarchical high-order neighbor prediction.
These LLM-based methods utilize LLMs to comprehend semantic information from heterogeneous textual attributes and generate unified textual descriptions or vector embeddings for downstream graph learning methods. The unified and high-quality outputs from LLMs greatly simplify the downstream learning process and enhance performance by addressing the challenges posed by cross-domain heterogeneity in textual attributes.
3.3.2 Multimodal Attributed Graph Learning
Compared to attributes within the same modality, multimodal attributes encompass information from various modalities, such as text, images, audio, and videos. These multimodal attributes offer a more comprehensive and enriched representation of the underlying data. By capturing different aspects of a node through diverse inputs from different sources, multimodal attributes enhance the contextual understanding and provide complementary information that single-modal data may lack [209, 210, 211]. However, this diversity also introduces additional challenges, as attribute heterogeneity across different domains manifests not only in data distribution but also in data formats. This variability complicates the unified processing of these attributes, requiring effective fusion methods to integrate information across different modalities [21, 212].
(1) Traditional Methods for Multimodal Attribute Modeling Traditional methods aim to learn alignment or fusion patterns across different modalities by training models to capture the underlying relationships between heterogeneous multimodal attributes. MMGCN [211] learns a multimodal graph convolution network on a user-item bipartite graph to learn modal-specific representations of users and micro-videos to better capture user preferences in different modalities. scMoGNN [213] utilizes different GNNs to learn representations of each cell-feature bipartite graph in different modalities, and finally concatenates these representations to fuse information from different modalities. [214] and [215] introduce an attention mechanism in graph neural networks to dynamically capture the importance of information from different modalities. To reduce the need for large amounts of labeled data and improve robustness, contrastive learning methods are proposed for modeling multimodal graph data. Joyful [216] designs a global contextual fusion module and a specific modalities fusion module to capture information in different scales, and then concatenates the representations from both modules to generate a unified representation vector for each node. These modules are trained by comparing positive and negative pairs in the corrupted graphs through edge perturbation and random masking. FormNetV2 [217] proposes a centralized multimodal graph contrastive learning strategy to learn fused representations from different modalities in one loss. HGraph-CL [218] introduces a hierarchical graph contrastive learning framework that builds intra-modal and inter-modal graphs, leveraging graph augmentations and contrastive learning to capture complex sentiment relations within and across different modalities.
While these traditional methods can align or fuse multimodal attributes in graph data, they still rely on abundant data to train the model and cannot easily generalize to an unseen domain. Compared to LLMs, these traditional methods are less adaptable to diverse data distributions and struggle to leverage pre-trained knowledge for more efficient generalization.
(2) LLM-based Methods for Multimodal Attribute Modeling Although LLMs are specifically designed for understanding natural languages and not naturally suited for handling multimodal attributes, they can be combined or aligned with models for other modalities. Leveraging their superior generalization ability, LLM-based methods eliminate the need for large amounts of application-specific training data, opening up new possibilities for addressing attribute heterogeneity across multimodal source domains.
Recent research on MLLMs focuses on developing advanced methods to align LLMs with models from other modalities [219, 220, 221, 222, 223]. Although these methods are not specifically designed for graph data, they can be used as powerful preprocessing tools for aligning multimodal attributes and generating unified representations. For example, by unifying textual and visual side information through a pretrained model Clip-ViT [224], LLMRec [167] effectively mitigates the multimodal heterogeneity and enhances the node features in the recommendation network, leading to significant performance improvements. MAGB [21] conducts experiments on a set of large multimodal attributed graph datasets, which demonstrates that MLLMs can effectively alleviate the biases from cross-domain multimodal heterogeneity. Recent approaches align encoders for different modalities by leveraging graph structures, making them more naturally suited for graph learning tasks. MMGL [173] utilizes LLMs and image encoders with adapter layers to embed text and image attributes, respectively. These embeddings are then combined with graph positional encodings to capture graph structure information and finally fed into LLMs to generate the corresponding outputs. GraphAdapter [174] introduces GNN-based adaptors for encoders of different modalities, which can better align these encoders based on the graph structure information. TouchUp-G [175] improves node features obtained from pretrained models of different modalities by adapting them to the graph structure, using a new metric called feature homophily to quantify the correlation between the graph and node features, which enhances GNN performance across different tasks and data modalities. Following this direction, UniGraph2 [176] leverages modality-specific encoders alongside a GNN and an MoE module to effectively unify multimodal features while preserving the underlying graph structure.
By aligning LLMs with models for different modalities, these methods can simultaneously understand the heterogeneity across modalities and map the multimodal attributes into a unified embedding space. These unified attribute representations simplify the downstream graph learning process and improve the performance of multi-modal graph learning tasks.
3.3.3 Structural Heterogeneous Graph Learning
Graph structures capture essential connectivity patterns that are crucial for real-world applications. However, graphs constructed from heterogeneous source domains can exhibit excessive heterogeneity in structural patterns. The structural heterogeneity stems from the inherent biases of distinct structural patterns across different source domains, which cannot generalize throughout the graph and may obscure truly valuable structural information. For instance, road network data collected from various cities always exhibit structural heterogeneity, where data from some cities may exhibit grid-like structural patterns and data from others have radial configurations [18]. Such heterogeneity hinders the models from capturing underlying generalizable structural features and therefore limits their performance on real-world applications like traffic flow prediction [225, 226] or route planning [227]. Traditional graph domain adaptation methods rely on training data from different domains and cannot generalize to unseen domains. In contrast, LLMs, with their superior semantic understanding and generalization abilities, can understand different graph structural patterns in a zero-shot manner [193], providing new opportunities for mitigating structural heterogeneity in different graph data from unseen domains.
(1) Traditional Methods for Heterogeneous Structure Modeling While many graph learning methods struggle to generalize across structures from different source domains with different data distributions, recent research investigates how to mitigate this issue by adapting a learned model from a source domain to a target domain. These graph adaptation methods provide solutions for eliminating the biases across different domains and help capture unified and generalizable embeddings for graph structures with excessive heterogeneity.
Earlier methods for graph domain adaptation, including DANE [20] and ACDNE [30], employ shared-weight GNNs to align the embedding spaces of different graphs and utilize a least squares generative adversarial network to regularize the distribution alignment, ensuring that the learned unified embeddings do not include domain-specific information. Different from these adversarial regularization methods, some approaches focus on directly aligning distributions from different domains using different metrics. For example, SR-GNN [23] addresses the domain shifts by regularizing the hidden layer distributions using central moment discrepancy. GDA-SpecReg [228] combines optimal transport theory and graph filter theory to derive a theoretical bound for graph domain adaptation. Based on this bound, the method utilizes Wasserstein-1 distance to regularize the node representation distributions. StruRW [31] proposes an effective approach to reduce conditional structural shifts by re-weighting the edges in the source graph. While most methods align distributions for spatial representations, DASGA [229] introduces a spectral-based approach by aligning the Fourier bases of the source and target graphs, which ensures that the label functions in both domains have similar coefficients in their respective bases.
While these traditional methods demonstrate superiority in mitigating cross-domain structural heterogeneity, they require separate source and target domain data for training, which limits their application in scenarios without such training data or with a large number of domains. Additionally, traditional GNNs typically rely on fixed k-hop neighborhoods, limiting their ability to generalize across highly diverse connectivity patterns from different source domains with varying hop distances. Furthermore, these methods are inherently dependent on training data, which constrains their ability to handle structures that deviate from the patterns observed during training [230].
Figure 11: Two different approachs for handling cross-domain structural heterogeneity with LLMs. (a) LLM-based methods convert graph structures into textual descriptions, enabling direct input to LLMs. (b) Hybrid methods align GNN-generated graph token embeddings with the native text token embeddings of LLMs in a unified representation space, thereby enhancing the structural comprehension capabilities of LLMs.
(2) LLM-based Methods for Heterogeneous Structure Modeling Different from traditional graph learning methods, LLM-based methods provide a more flexible alternative by representing graph structures through unified text-based descriptions, which are not restricted by a fixed k-hop structural neighborhood assumption. Instead of relying on predefined graph processing mechanisms, LLMs, with superior comprehension and generalization abilities, can understand unseen structural patterns in different domains. Since LLMs are designed to process natural language inputs, a direct way to bridge the gap between graph structures and the input format of LLMs is to transform structural information into textual descriptions, as depicted in Figure 11(a). Here we survey recent advances in structure-to-text transformation methods that enable LLMs to understand heterogeneous structural patterns. Although some methods are not specifically designed for mitigating cross-domain structural heterogeneity, they provide strong potential for addressing the associated challenges.
Recent works have explored how to effectively transform graph structures into text to enable better understanding and reasoning by LLMs. Early methods [177, 178, 179, 180] experimented with existing graph textualization techniques, such as structure description (like node sequences and edge sequences) and formal languages (like graph markup language [231]). While LLMs can comprehend graph structures from these natural language descriptions, their results demonstrate that the choice of different textualization techniques can significantly affect the performance of LLMs on different graph tasks. Different from previous methods that only consider natural language format for representing graph structures, GITA [181] transforms graph structures into both text and images to obtain a better understanding of structural information from different modalities, which are then fed into a Vision-Language Model (VLM) for addressing graph tasks. LLM4Hypergraph [182] designs low-order and high-order structure languages to transform hypergraph structures into natural languages. These studies have significantly advanced the exploration of structure-to-text approaches for LLMs to comprehend graph topological structures.
Beyond merely understanding topological structures, textual descriptions can also incorporate the attributes of nodes and edges, providing a richer context that enables LLMs to capture the intricate features of the graph structure. GraphText [183] constructs a graph-syntax tree to preserve the hierarchical structure and traverses the tree structure to generate a graph text sequence, where the node attributes are incorporated in the description of leaf nodes. WalkLM [184] generates paths by attributed random walks and textualizes the paths into natural languages for processing in LLMs. To effectively capture cross-group connections while minimizing noisy nodes, Path-LLM [185] employs shortest paths to generate structural sequences. InstructGLM [186] describes graph structures using neighbors in different scales and utilizes instruction tuning to finetune an LLM to better perform graph tasks. The textual node attributes are concatenated with the node index to better incorporate semantic information. MuseGraph [187] incorporates neighbors and paths together to textualize graph structures, which can capture both local connectivity and complex path-based relationships between nodes. Graph-LLM [188] transform text-attributed graphs into natural languages and demonstrates that LLMs can understand the graph structures from language descriptions and demonstrate remarkable zero-shot performance on graph tasks.
By converting diverse graph structures into natural language descriptions, these structure-to-text transformation methods enable LLMs to interpret and reason across varying connectivity patterns in a unified manner. These approaches allow LLMs to leverage their pre-trained linguistic knowledge to identify valuable relational patterns and facilitate more flexible reasoning across different graph topologies with excessive heterogeneity.
(3) Hybrid Methods for Heterogeneous Structure Modeling While structure-to-text transformation methods offer a straightforward way to bridge the gap between graph structures and the input format of LLMs, they heavily depend on the language comprehension ability of LLMs to correctly interpret the input graph structures. Some recent works explore learning explicit structure representations using traditional graph learning methods like GNNs and aligning these representations with the token space of LLMs, as depicted in Figure 11(b). These hybrid methods leverage the structural modeling capabilities of traditional graph learning methods alongside the comprehension and reasoning strengths of LLMs for more effective graph structure modeling.
GNP [189] employs a GNN to encode the knowledge graph structure into graph tokens and subsequently uses projectors to map these tokens into the text token space of an LLM. Since the projected graph tokens share the same embedding space as the original text tokens of the LLMs, they can be seamlessly concatenated with language instructions and fed into the LLM. This integration allows the LLM to leverage both graph structures and textual context to tackle various graph-related tasks. GraphToken [190] generates graph tokens through a GNN for general graphs and aligns the graph tokens with the token space of a frozen LLM. GraphGPT [191] introduces a dual-stage instruction tuning framework that directly aligns graph tokens with LLMs using both self-supervised and task-specific instruction tuning. LLaGA [192] reorganizes graph nodes into structure-aware sequences based on two different templates, then maps these sequences into the token embedding space of LLMs through a versatile projector, enabling the model to handle various graph tasks while maintaining the general-purpose nature of LLMs. TEA-GLM [193] utilizes feature-wise contrastive learning with token embeddings, where principal component analysis is applied to the token embeddings of an LLM to obtain principal components that serve as coordinate axes. These axes are used to map node representations generated from a GNN into the token space of the LLM. To generate graph tokens for heterogeneous graphs, HiGPT [194] introduces an in-context heterogeneous graph tokenizer that encodes diverse node and edge types by using a language-based parameterized heterogeneity projector, dynamically generating graph tokens that represent the semantic relationships in heterogeneous graphs.
These hybrid methods incorporate structure representations from traditional graph learning methods to facilitate the structure comprehension ability of LLMs, providing more effective solutions for mitigating the challenges from structural heterogeneity. Additionally, the use of compact graph tokens reduces input size and enables LLMs to process larger structures within their context window, which is crucial for understanding heterogeneous structural patterns across varying scales.
3.3.4 Summary of Heterogeneity
LLMs, pre-trained on diverse datasets, are inherently capable of integrating information from various sources into a unified semantic space. The reviewed studies show that LLM-integrated models can transfer knowledge across different graph domains without the need for domain-specific retraining. By leveraging the generalizability of LLMs, these models are able to mitigate different forms of cross-domain heterogeneity, including textual attribute heterogeneity, multimodal attribute heterogeneity, and structural heterogeneity. As a result, they demonstrate superior performance in addressing cross-domain heterogeneity, particularly in scenarios where traditional methods are ineffective or inapplicable due to the lack of sufficient training data. Experiments in these papers show improved performance on cross-domain node classification, link prediction, and graph classification tasks when LLMs are used to either encode heterogeneous content or assist in aligning representations.
3.4 Dynamic Instability in Graphs
Figure 12: Illustration of dynamic instability in graphs. The figure presents three categories of dynamic instability over two timestamps ($t_{1}$rightarrow $t_{2}$): (1) Node Feature Dynamics, where node attributes change over time; (2) Structural Dynamics, showing newly added node (solid dark purple circle) and newly removed node (gray dashed circle); (3) New Node/Edge Types, depicting the emergence of entirely new node and relationship types, represented by a solid dark purple triangle and dashed edges.
Many real-world systems, from social networks to knowledge bases, are best represented as graphs whose structure and attributes change over time (see Figure  12). This constant evolution implies a shifting distribution, where patterns observed in the past may not hold true in the future. While gradual evolution poses modeling challenges, the situation becomes significant when the graph undergoes rapid, large-scale, or highly unpredictable changes that alter the graph’s characteristics and create significant challenges for model training and inference [232].
Dynamic graphs’ instability is challenging to model due to their temporal and structural variability [233]. In static settings, models learn stable representations from collected graph data. In contrast, dynamic graphs experience frequent changes in nodes, edges, and attributes. Models struggle to generalize without adaptive mechanisms when faced with structural shifts or abrupt transformations [234]. Before the introduction of LLMs, researchers addressed dynamic graphs using snapshot-based and incremental methods. These approaches discretized the graph into time-specific snapshots for independent training [32], or updated models as new nodes and edges appeared [26]. While effective in some settings, such methods often rely on localized updates and struggle to capture long-range temporal dependencies, particularly in large-scale or rapidly changing graphs.
LLMs have recently emerged as a powerful new tool with the potential to address some challenges in dynamic graph learning (Figure 13). By leveraging their strengths in natural language understanding, sequential data processing, few-shot learning, and complex reasoning, LLMs offer novel ways to interpret and model graph dynamics. They can process textual information associated with nodes or edges and understand the semantic context driving structural changes. This infusion of semantic reasoning capabilities opens promising research routes for creating more robust and adaptive models capable of navigating the complexities of dynamic instability. The relevant references and categorization are presented in Table 4.
Table 4: Literature Summary of LLM-based Methods for Handling Dynamic Instability in Graphs.
Domain Type Reference (Method—Citation) Querying and Reasoning Forecasting & Reasoning ICL[235], zrLLM[236], CoH[237], TG-LLM[238], LLM4DyG[239] QA & Interpretability TimeR4[240], GenTKGQA[241], Unveiling LLMs[242] Generating and Updating Generating Structures FinDKG[243], GenTKG[244], Up To Date[245], PPT[246], LLM-DA[247] Generating Insights & Representations TimeLlama[248], RealTCD[249], DynLLM[250] Evaluation and Application Model Evaluation Dynamic-TempLAMA[251], DARG[252] Downstream Applications AnomalyLLM[83], MoMa-LLM[253], TRR[254]
Figure 13: A unified framework integrating LLMs for addressing dynamic instability in graphs. Dynamic changes in node features, structure, and new node or edge types are captured by temporal graph embeddings. These dynamic embeddings, combined with context (triples and historical facts), serve as inputs to the LLM. The LLM then performs reasoning tasks and outputs the final prediction or response.
3.4.1 Querying and Reasoning in Dynamic Graphs
This category includes work that uses LLMs to query, reason about, or retrieve knowledge from dynamic graphs. These approaches analyze or validate existing graph structures rather than actively generating or modifying node and edge features. LLMs act as information retrieval or reasoning tools, querying, analyzing, or inferring over existing graph structures and knowledge to discover new relationships or validate known ones. The key aspect is that LLMs assist in analyzing and understanding the graph’s dynamic evolution.
(1) Traditional Methods Before the emergence of LLMs, querying and reasoning on dynamic graphs relied on specialized graph representation learning and temporal modeling methods. For example, Temporal Graph Networks (TGNs) [233], and related methods use memory modules combined with graph neural networks to process continuous-time dynamic graphs and capture node evolution. Methods like DyRep [33] and TGAT [255] focus on learning node representations that reflect temporal changes in neighborhood structures and interaction patterns. For reasoning on Temporal Knowledge Graphs (TKGs), such as link prediction, traditional approaches employed embedding techniques or rule-based systems to capture changes in entities and relations over time [256]. While effective for specific tasks, these methods struggle to fully use rich semantic information and face challenges in handling zero-shot relations or performing complex multi-step reasoning [257].
(2) Forecasting and Reasoning Current research now explores the semantic understanding and reasoning capabilities of LLMs to address querying and reasoning challenges in dynamic graphs, particularly for TKG forecasting and reasoning. One surprising finding is that LLMs, even without fine-tuning, can achieve performance comparable to specialized TKG models on forecasting tasks simply by using In-Context Learning (ICL) with historical facts converted to text. Their performance holds even when entity names are replaced with numerical IDs, suggesting LLMs can leverage structural and temporal patterns in the context [235]. To deepen LLM reasoning, CoH (Chain-of-History) [237] proposes a method to explore high-order historical information step-by-step, overcoming the limitation of relying only on first-order history and improving temporal reasoning, especially as a plug-in module for graph-based models. TG-LLM [238] trains an LLM to translate text context into a latent temporal graph and then uses Chain-of-Thought (CoT) reasoning over this graph, enhancing generalizable temporal reasoning. To address the challenge of unseen relations in TKGs, zrLLM [236] uses LLMs to generate semantic representations from relation descriptions, enabling embedding models to recognize zero-shot relations via semantic similarity. Furthermore, LLM4DyG [239] introduces a systematic evaluation of LLMs’ spatial-temporal understanding on general dynamic graphs. It introduces a benchmark and proposes the “Disentangled Spatial Temporal Thoughts” prompting method to improve performance on tasks like link prediction and node classification, although challenges remain with large or dense dynamic graphs.
(3) QA and Interpretability Another line of research focuses on using LLMs for more complex information retrieval from dynamic graphs, such as Temporal Knowledge Graph Question Answering (TKGQA), and understanding the LLMs’ own reasoning processes. TKGQA requires understanding complex temporal constraints in questions and retrieving answers from dynamic knowledge, a task where traditional methods struggle with semantics. TimeR4 [240] introduces a time-aware Retrieve-Rewrite-Retrieve-Rerank framework. It uses LLMs and retrieved TKG facts to handle time constraints and reduce temporal hallucination by rewriting questions and reranking retrieved facts. Similarly, GenTKGQA [241] employs a two-stage approach: the LLM first mines constraints to guide subgraph retrieval, then generates answers by fusing GNN signals with text representations via instruction tuning. These methods demonstrate the potential of retrieval-augmented LLMs for complex dynamic knowledge QA. Additionally, to understand how LLMs process factual knowledge, Unveiling LLMs [242] uses dynamic knowledge graphs as a tool to interpret LLM reasoning. It decodes internal token representations layer-wise during fact verification, revealing how factual representations evolve within the LLM.
3.4.2 Generating and Updating in Dynamic Graphs
This category includes papers that use LLMs to actively generate new nodes, edges, or their features or to update dynamic graph structures. These approaches typically serve to complete, reconstruct, or adapt the graph structure to dynamic changes. LLMs function as generators or updaters, creating new node/edge attributes or descriptions to fill in missing information or updating the graph structure in real time based on changes. The focus is on utilizing LLMs’ generative capabilities to actively create or modify the graph structure to cope with dynamics.
(1) Traditional Methods Before LLMs became widespread, addressing dynamic graph generation and updates relied mainly on graph generative models and dynamic embedding techniques. Models like GraphRNN [258] and GRAN [259] focused on generating static graph structures; extending them to dynamic settings proved challenging. For updating representations in dynamic graphs, methods like DyGEM [260] adapted node embeddings at each time step to reflect graph changes. Knowledge graph updates often depended on manual editing, rule-based systems, or specific database maintenance procedures, which were difficult to automate and could not respond rapidly to real-world changes. These traditional methods generally lacked the ability to use external unstructured information (like text) to guide updates and had limitations in generating semantically consistent and structurally complex graph evolution patterns.
(2) Generating Structures Researchers now use LLMs’ generative power to create or maintain evolving knowledge structures. The FinDKG [243] demonstrates the potential of LLMs as dynamic knowledge graph (DKG) generators; their ICKG model builds a DKG directly from financial news text to capture market trends. To address the issue of outdated information in knowledge graphs, Up To Date [245] proposes a method combining LLM reasoning and RAG. It automatically identifies potentially outdated facts in a KG and retrieves information from trusted sources to generate accurate corrections, enabling automated KG maintenance. For TKG completion and forecasting, PPT [246] converts TKG facts and time intervals into prompted natural language sequences, using a masked language modeling task for completion. GenTKG [244] also adopts a generative forecasting approach for TKGs. It uses a retrieval-augmented framework with parameter-efficient instruction tuning to generate future facts, achieving strong performance and generalization even with minimal training data. Furthermore, LLM-DA [247] innovatively uses an LLM to generate temporal logical rules from historical TKG data to guide reasoning. A dynamic adaptation strategy updates these rules based on new events, allowing the model to adapt to knowledge evolution without retraining the LLM.
(3) Generating Insights and Representations Beyond directly generating graph structures or facts, LLMs generate higher-level information like explanations, causal hypotheses, or enhanced node representations to aid dynamic graph analysis and applications. TimeLlama [248] focuses on explainable temporal reasoning. It not only predicts future events but also uses an LLM to generate natural language explanations based on historical TKG paths, increasing model trustworthiness. To explore the underlying mechanisms of dynamic systems, RealTCD [249] employs LLMs to process textual information (e.g., system logs) and integrate domain knowledge. Through LLM-guided meta-initialization, it improves the quality of temporal causal discovery, especially in industrial scenarios lacking intervention targets. For dynamic recommendation systems, DynLLM [250] uses LLMs to generate multi-dimensional user profiles (e.g., interests, preferred brands) and their embeddings from the textual features of purchase histories. These LLM-generated, dynamically updated user profiles enrich user node information, improving the recommendation system’s ability to adapt to changing user preferences. These studies show that LLMs can produce basic graph elements and high-level semantic information, offering new ways to understand and model the complex evolution of dynamic graphs.
3.4.3 Evaluation and Application in Dynamic Graphs
This category includes work that uses LLMs to evaluate the effectiveness of dynamic graph learning models or apply LLMs in downstream tasks such as link prediction, node classification, or recommendation systems. LLMs serve as evaluation tools or modules within downstream applications, assessing the performance of dynamic graph models or providing additional guidance during training or inference. The emphasis is on using LLMs’ comparative, assessment, or decision-support capabilities to handle tasks in dynamic environments.
(1) Traditional Methods Before using LLMs in these tasks, researchers relied on models that predicted future links or node states based on historical graph data [261]. However, standard evaluation metrics often fail to capture the predictions’ semantic quality or real-world plausibility [262]. These metrics also struggled to assess how models handled unexpected structural changes or evolving concepts. In downstream applications, tasks such as dynamic anomaly detection typically use statistical methods to identify shifts in graph structure or connectivity patterns [263]. Dynamic recommendation systems often relied on sequential models that analyzed user interaction histories [264] or time-aware collaborative filtering. Although these approaches were effective for specific problems, they lacked the ability to perform deeper semantic reasoning.
(2) Model Evaluation LLMs provide new perspectives for evaluating model performance and robustness in dynamic settings. Dynamic-TempLAMA [251] presents a dynamic benchmarking framework specifically designed to assess how well pre-trained language models (MLMs) handle temporal concept drift—the evolution of factual knowledge over time. The framework dynamically creates time-sensitive test sets from Wikidata and evaluates MLMs through multiple views (probing, generation, scoring) to determine if their internal knowledge is outdated. DARG [252] addresses the limitations of static benchmarks by proposing a method to dynamically generate new evaluation data. It introduces changes to the reasoning graphs of existing benchmark samples to create novel test data with controlled complexity and diversity, using a code-augmented LLM to ensure label correctness. This enables adaptive evaluation of LLMs’ reasoning capabilities as they evolve. Both works highlight the importance of moving beyond static, snapshot-based evaluations toward more dynamic, adaptive approaches, especially as both models and world knowledge constantly change.
(3) Downstream Applications LLMs are also embedded directly into downstream applications that process dynamic graph data, serving as core reasoning or decision-making components. AnomalyLLM [83] uses LLM knowledge for few-shot anomaly edge detection in dynamic graphs. By aligning edge representations with word embedding prototypes and using in-context learning, the method effectively identifies emerging anomaly types with few labeled examples, demonstrating LLMs’ potential for adapting to changing environments like cybersecurity. In robotics, MoMa-LLM [253] integrates an LLM with dynamically updated, open-vocabulary scene graphs representing an explored environment. The LLM reasons over this evolving graph to guide a mobile robot in long-horizon, interactive object search tasks, showcasing how LLMs can integrate dynamic spatial-semantic information for high-level planning. The financial sector also leverages LLMs; TRR (Temporal Relational Reasoning) [254] uses an LLM-based framework mimicking human cognition (memory, attention) to detect potential stock portfolio crashes. It reasons over dynamically generated temporal relational information extracted from news to assess the aggregated impact of evolving events, which is useful for rare events lacking historical data. These applications demonstrate LLMs acting as powerful reasoning engines in complex real-world tasks that require understanding and responding to dynamically changing structured information.
3.4.4 Summary of Dynamic Instability
The integration of LLMs with dynamic graph learning presents a promising direction for addressing the inherent challenges posed by evolving graph structures. Leveraging their powerful sequence modeling capabilities and the flexibility of natural language encoding, LLMs can effectively adapt to distribution shifts and temporal variations that traditional static graph models struggle to capture. Moreover, LLMs can incorporate external temporal knowledge into the graph reasoning process, enhancing predictive power and enabling anticipation of future changes rather than merely reacting to them. Early empirical evidence suggests that LLM-augmented dynamic graph models exhibit greater robustness and sustained performance over time. Nevertheless, these benefits come with increased computational costs and complexity, requiring careful architectural design and error mitigation strategies to prevent error propagation and maintain reliability in long-term deployment.
4 Future Directions and Open Challenges
4.1 Incompleteness in Graphs
4.1.1 Controlling Hallucination in LLM-based Graph Completion
LLMs hold strong potential for completing missing nodes, edges, or attributes in graphs by generating latent knowledge from textual context. However, a key challenge lies in mitigating hallucination—the generation of plausible but incorrect or structurally inconsistent information. Ensuring that LLM-generated content aligns with the graph’s schema and topology requires future work on structure-constrained decoding, confidence-aware integration, and post-generation validation using GNNs or external knowledge bases. Jointly training LLMs with graph-aware objectives may further reduce hallucination and improve the structural fidelity of completions.
4.1.2 LLMs’ Robust Inference on Incomplete Data
LLMs demonstrate robustness when faced with missing or noisy inputs, but their reasoning remains opaque and uncontrolled. Future research should investigate how to guide LLMs toward faithful, graph-consistent inference, particularly when data is partially observed. One avenue is to design structure-aware prompts or adapters that encode local subgraph structure, enabling LLMs to learn inductive biases from graph topology rather than relying solely on pretraining knowledge. Additionally, LLMs could be enhanced to support iterative reasoning over subgraphs, where initial predictions are refined through multiple interaction rounds with the graph context. This would mimic how GNNs aggregate information layer by layer, and could bridge the reasoning styles of symbolic and connectionist methods.
4.1.3 Benchmarks and Interpretability in LLM-based Graph Learning
Despite growing interest, there is a lack of standardized benchmarks and evaluation protocols tailored for LLM-based graph learning. Future work should establish datasets and tasks that reflect real-world challenges such as incomplete graphs, noisy node attributes, and open-world reasoning. Furthermore, interpretability remains a major hurdle: how can we trust LLM-driven graph predictions? Tools that visualize attention over graph structures, track reasoning chains within LLMs, or contrast outputs under counterfactual graph edits could greatly enhance model transparency.
4.2 Imbalance in Graphs
4.2.1 Semantic-Topology Joint Debias
To address the imbalance between semantic logic and topological patterns in graph data, the focus of research is on constructing a text-graph joint representation space to uncover the deep interconnections between multimodal information. Leveraging the causal reasoning ability of LLMs, it is possible to identify the semantic roots of structural imbalance, such as topological sparsity caused by the logical absence of long-tail relationships in knowledge graphs. By generating logically constrained virtual edges or subgraph structures, this approach can reconstruct the global distribution of the graph from a causal perspective, achieving a synergistic enhancement of semantic coherence and topological integrity. Typical application scenarios include: in knowledge graph completion tasks, LLMs can be used to parse the implicit semantic relationships between entities and generate virtual triples; in social network analysis, user behavior text can be combined to infer potential social links, correcting network structural flaws caused by data collection biases.
4.2.2 Dynamic Cross-Modal Transfer
To mitigate class or structural bias in graph learning tasks caused by data sparsity, lightweight adaptation techniques need to be developed for the dynamic transfer of LLMs’ open-domain knowledge to graph structures. By designing cross-modal projection layers and attention alignment mechanisms, LLMs’ semantic representations can be mapped to a unified vector space with graph topological features. In molecular property prediction scenarios, this can achieve spatial matching between atomic-level text descriptions and molecular graphs, or in social network analysis, it can integrate user profile text with community structure features. This technology breaks through the closed training approach of traditional graph models, utilizing LLMs’ generalized knowledge to compensate for deficiencies in domain-specific data.
4.2.3 Trustworthy Generation & Efficiency Optimization
To ensure semantic validity when LLMs generate minority-class nodes or edges in imbalanced graphs, future work should focus on suppressing hallucinations through structure-aware verification. By cross-validating GNN-derived topological patterns with LLM-generated semantics, models can dynamically filter or penalize low-confidence, type-inconsistent outputs. Prototype alignment and discriminative validators further help constrain generation within class-specific semantics. Meanwhile, parameter-efficient tuning (e.g., LoRA) and multimodal pretraining enable scalable deployment and real-time adaptation to long-tail patterns in dynamic graphs, such as correcting abnormal transaction structures via time-aligned textual streams.
4.3 Cross-Domain Heterogeneity in Graphs
4.3.1 Context-Aware Heterogeneous Attribute Learning
Existing LLM-based methods for handling attribute heterogeneity typically treat LLMs or MLLMs as powerful preprocessing tools for generating unified embeddings for node attributes. However, modeling attributes in isolation without leveraging the abundant contextual information leads to suboptimal representations. For instance, the semantic meaning of a textual node attribute can be enriched by incorporating other textual attributes or even multimodal attributes associated with the same node. Additionally, the connectivity patterns of a node offer important relational information that enhance the interpretation of its attributes. Moreover, different tasks emphasize different aspects of attributes, making task-related context essential for generating embeddings aligned with specific learning objectives. Future research can explore integrating semantic, structural, and task-related context information to enhance heterogeneous attribute learning.
4.3.2 Symmetry-Preserving Heterogeneous Structure Learning
LLMs exhibit strong capabilities in understanding and reasoning over heterogeneous graph structures represented in natural language. However, they do not explicitly incorporate graph symmetries such as invariance. For instance, the representation of a graph should remain consistent regardless of the ordering of its input node sequence, as different orderings do not alter the underlying structure. Existing LLM-based methods fail to fully exploit such graph symmetry, making it difficult to determine whether variations in the input graph descriptions stem from distinct heterogeneous structures from different source domains or merely the same structure under different permutation actions. To address this issue, one possible approach is to design prompts that explicitly describe the relevant symmetries and provide examples in the context, helping LLMs understand the symmetric constraints of the input graph. Additionally, LLMs can be fine-tuned using contrastive learning to distinguish whether two differently ordered node sequences represent the same underlying graph structure to capture the symmetry knowledge.
4.3.3 Effective Integration of GNNs and LLMs
Although LLMs excel at capturing semantic and structural knowledge from graph data represented in text, they cannot utilize the entire graph data due to limited context window sizes. In contrast, GNNs can be easily trained on large graphs due to the efficient message passing mechanism. Current methods for handling cross-domain heterogeneity explore using LLMs as attribute encoders for GNNs or generating graph tokens with GNNs for LLM reasoning. These integration methods are limited by the expressiveness of GNNs, which may fail to detect some critical substructures [265, 266]. To better utilize the advantages of both methods, one potential direction is to design a dual architecture, where the final prediction could be based on the outputs of both LLMs and GNNs. Another direction is to distill the knowledge from a GNN trained on the entire graph into an LLM, allowing the LLM to effectively capture the knowledge from the graph data without finetuning on the whole graph.
4.4 Dynamic Instability in Graphs
4.4.1 Querying and Reasoning
A key challenge is scaling LLM-based reasoning to large and high-frequency dynamic graphs while keeping inferences grounded in the underlying graph structure. Current models often struggle with hallucinations and fail to reflect the evolving nature of real-world graphs. Another limitation is the difficulty in representing complex temporal dependencies in a form that LLMs can effectively process. Future work should focus on building tighter integrations between LLMs and GNNs, developing pretraining strategies that incorporate temporal data, and designing efficient, time-aware prompting and retrieval methods. Improving the ability of LLMs to perform causal reasoning and making their decision processes more transparent are also important directions for research.
4.4.2 Generating and Updating
Current LLMs face challenges in generating graph elements that are faithful to known facts, structurally consistent, and aligned with domain constraints. Frequent updates to dynamic graphs also raise concerns about computational cost and system stability. Controlling the generation process remains challenging, especially when outputs need to be incorporated into existing graph structures. Future research should investigate constrained generation techniques, efficient methods for incremental updates, and mechanisms for integrating feedback from downstream tasks. Extending LLMs to model more complex graph dynamics and to incorporate signals from multiple modalities may enhance their ability to support real-world applications.
4.4.3 Evaluation and Application
LLM-based evaluation methods must address issues of reliability, bias, and limited generalization. Static benchmarks fail to reflect the changing nature of dynamic graphs and the models that operate on them. The high cost and latency of LLMs also present barriers to real-time deployment. Future directions include developing adaptive evaluation frameworks that combine LLM-based reasoning with traditional performance metrics and exploring human-in-the-loop strategies to improve evaluation quality. Research should also focus on building lightweight deployment architectures and designing LLM-based systems that adapt proactively to changing graph environments while carefully addressing fairness and ethical concerns.
5 Conclusion
In this survey, we present a comprehensive review of how recent advances in large language models can address four fundamental challenges in graph learning that arise from mismatches between ideal assumptions and real-world graphs. These core challenges are (1) Incompleteness of graph structures or attributes, (2) severe Imbalance in node/edge distributions, (3) Cross-Domain Heterogeneity in semantics and structure, and (4) Dynamic Instability due to evolving topologies and interactions. Each of these issues highlights a facet of graph data that can confound traditional graph learning methods and limit their generalization ability.
We reviewed a wide array of LLM-integrated approaches proposed to tackle these challenges, explaining them through a unifying conceptual framework. The findings across the literature demonstrate that LLMs offer significant advantages when combined with graph learning techniques. For incomplete graphs, LLMs provide powerful semantic reasoning and external knowledge to infer missing node attributes and relationships, effectively acting as intelligent imputers to complete graphs beyond what purely graph-based methods can do. In imbalanced graphs, LLMs can generate synthetic data and augment feature spaces to alleviate class or structural imbalance, introducing new examples and nuanced features that help models learn minority patterns and thus improve robustness against bias. For heterogeneous graphs, LLMs produce unified semantic embeddings and serve as a bridge across modalities and domains, harmonizing disparate attribute types and structural schemas into a common representation. In dynamic graphs, LLMs enable advanced temporal reasoning and contextual analysis, from anticipating future links to explaining changes, thereby helping graph models continuously adapt over time.
Looking ahead, several open questions and promising directions emerge. First, improving the efficiency of LLM-graph integration remains critical, particularly for large-scale or real-time applications. Second, enhancing the explainability of LLM-driven graph reasoning will foster trust in high-stakes domains like healthcare and finance. Third, developing frameworks for continuous adaptation to dynamic graphs—without catastrophic forgetting—requires further exploration. Fourth, bridging the gap between textual knowledge encoded in LLMs and structural graph signals could unlock novel cross-modal learning approaches. Finally, addressing the issue of hallucination in LLM-based graph generation is essential to ensure that generated content is both semantically accurate and structurally consistent, fostering more reliable and trustworthy graph learning.
In closing, the integration of large language models with graph learning represents a new frontier that brings together the best of two worlds: the deep semantic understanding from NLP and the structured relational modeling from graph ML. The synergy between these fields is already yielding models that are more robust, knowledgeable, and adaptable than those based on either technology alone. As research continues along the directions identified, we anticipate rapid progress toward graph learning systems with far greater generality and intelligence, capable of reasoning about the rich, dynamic networks that underlie real-world data.
Author Contributions
Mengran Li: Conceptualization, Methodology, Formal analysis, Writing – original draft, Visualization. Pengyu Zhang: Conceptualization, Methodology, Formal analysis, Writing – original draft, Visualization. Wenbin Xing: Conceptualization, Data curation, Investigation, Writing – original draft, Visualization. Yijia Zheng: Conceptualization, Data curation, Investigation, Writing – original draft, Visualization. Klim Zaporojets: Software, Validation, Writing – review & editing. Junzhou Chen: Supervision, Writing – review & editing. Ronghui Zhang: Supervision, Project administration, Funding acquisition, Writing – review & editing. Yong Zhang: Supervision, Resources. Siyuan Gong: Supervision. Jia Hu: Supervision, Resources. Xiaolei Ma: Supervision, Resources. Zhiyuan Liu: Supervision, Writing – review & editing. Paul Groth: Supervision, Writing – review & editing. Marcel Worring: Supervision, Writing – review & editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
This work was supported in part by the Shenzhen Science and Technology Program under Grant JCYJ20240813151129038; in part by the National Natural Science Foundation of China under Grant 52172350, and Grant 51775565; in part by the Guangdong Basic and Applied Research Foundation under Grant 2022B1515120072; and in part by the Science and Technology Planning Project of Guangdong Province under Grant 2023B1212060029.